OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary. a = ?, b = 5, c = 8

OpenStudy (anonymous):

which sides of the triangles correspond to a, b, c?

OpenStudy (amistre64):

and how are we to know the pythag will even work?

OpenStudy (anonymous):

Yeah, how will we know that this is a right triangle?

OpenStudy (anonymous):

The equation for the Pythagorean Theorem is a^2+b^2=c^2 so just input the numbers a^2+25=64 a^2+25-25=64-25 subtract 25 from both sides a^2=39 square root both sides a=6.24499 round to the nearest tenth it should equal 6.2, but I would have to agree with the others somewhat make sure you state whether the triangle is a right triangle or not as the Pythagorean Theorem doesn't work with any other type of triangle

OpenStudy (anonymous):

a = √ (8² - 5²) = √39 = 6.2